# Bayesian WeakS-to-Strong from Text Classification to Generation

### v2

@@ Abstract
Advances in large language models raise the question of how alignment techniques will adapt as models become increasingly complex and humans will only be able to supervise them weakly. Weak-to-Strong mimics such a scenario where weak model supervision attempts to harness the full capabilities of a much stronger model. This work extends Weak-to-Strong to WeakS-to-Strong by exploring an ensemble of weak models which simulate the variability in human opinions. Confidence scores are estimated using a Bayesian approach to guide the WeakS-to-Strong generalization. Furthermore, we extend the application of WeakS-to-Strong from text classification tasks to text generation tasks where more advanced strategies are investigated for supervision. Moreover, direct preference optimization is applied to advance the student model’s preference learning, beyond the basic learning framework of teacher forcing. Results demonstrate the effectiveness of the proposed approach for the reliability of a strong student model, showing potential for superalignment.  [^1] Code will be available upon acceptance.
@@ 3.3 Bayesian WeakS-to-Strong
Consider a weak label from the $i\text{th}$ weak model $\bm{y}_{w}^{(i)}$, which is a one-hot vector with $y_{w}^{(i,k)}$ being one if it belongs to class $k$ and zero otherwise. $\bm{y}_{w}^{(i)}$ is sampled from a categorical distribution of weak labels $\operatorname{Cat}(\bm{\pi})$, where each component ${\pi}_{k}$ corresponds to the probability assignment over the possible classes $\bm{y}_{w}^{(i)}\sim\mathrm{P}(\bm{y}|\bm{\pi})=\operatorname{Cat}(\bm{\pi}).$ EDL places a Dirichlet prior over the categorical distribution representing the probability of each categorical probability assignment, hence modelling second-order probability $\bm{\pi}\sim\mathrm{p}(\bm{\pi}|\bm{\alpha})$, where $\bm{\alpha}$ is the hyperparameter of the Dirichlet prior. The strong model $\bm{f_{\Lambda}}$ is trained to predict $\bm{\alpha}$ for each input by minimizing the negative log-likelihood of sampling $\bm{y}_{w}^{(i)}$ given the predicted Dirichlet prior:
@@
where $K$ is the number of classes, $\alpha_{0}=\sum_{k=1}^{K}\alpha_{k}$ is the Dirichlet strength, and $y_{w}^{(i,k)}$ is the $k^{th}$ value of label $\bm{y}_{w}^{(i)}$. Following Sensoy et al. (2018), a regularization term $\mathcal{L}_{\text{REG}}^{(i)}$ (see Appendix G for details) is added to calibrate uncertainty estimation, resulting in the EDL loss $\mathcal{L}_{\text{EDL}}^{(i)}=\mathcal{L}_{\text{NLL}}^{(i)}+\lambda_{\text{EDL}}\mathcal{L}_{\text{REG}}^{(i)}$ where $\lambda_{\text{EDL}}$ is the coefficient.
@@
$\mathcal{L}_{\text{EDL}}(\bm{f_{\Lambda}}(\bm{x}),\{\bm{y}_{w}^{(i)}\}_{i=1}^{N})=\sum_{i=1}^{N}\lambda_{i}\sum_{k=1}^{K}p_{k}\mathcal{L}_{\text{EDL}}^{(i)}(\bm{f_{\Lambda}}(\bm{x}),\hat{\bm{y}}_{w}^{(i)}))$ (4)
where $\hat{\bm{y}}_{w}^{(i)}$ is the predicted class, and $\lambda_{i}$s are hyperparameters set to the same values as used in the Naive Multi-Weak approach. As a result, the auxiliary confidence loss described in Eqn. (1) is adapted for Bayesian WeakS-to-Strong as follows:
@@ 4.1 Probability estimation for weak sequence labels
$\mathrm{P}(s_{1})=\frac{e^{C_{s}(s_{1})}}{e^{C_{s}(s_{1})}+e^{C_{s}(s_{2})}}\mathrm{P}(W),\qquad\mathrm{P}(s_{2})=\frac{e^{C_{s}(s_{2})}}{e^{C_{s}(s_{1})}+e^{C_{s}(s_{2})}}\mathrm{P}(W),$ (7)
@@ 4.2 DPO for sequence generation optimization
where $\bm{y}_{s}(m)$ is the $m$th output sequence generated by the strong model, $\mathrm{P}(\bm{y}_{s}(m)|\bm{\theta}_{i})$ is computed using the $i$th weak model with model parameters $\bm{\theta}_{i}$, and $\lambda_{i}$ is the weight assigned to the $i$th weak model. The dispreferred sequence can be computed similarly by
$\bm{y}_{r}=\operatorname*{arg\,min}_{m,m=1,2,\dots,M}\mathrm{P}(\bm{y}_{s}(m))=\operatorname*{arg\,min}_{m,m=1,2,\dots,M}\sum_{i=1}^{N}\lambda_{i}\mathrm{P}(\bm{y}_{s}(m)|\bm{\theta}_{i}).$ (9)
@@ 5.2 Models and baselines
Models. For classification, the Qwen-7B (Bai et al., 2023) [^2] https://huggingface.co/Qwen/Qwen-7B model was applied as the strong student model. Five models were used as weak teachers: GPT2-Large (Radford et al., 2019) [^3] https://huggingface.co/openai-community/gpt2-large, OPT-1.3B (Zhang et al., 2022) [^4] https://huggingface.co/facebook/opt-1.3b, Pythia-1.4B (Biderman et al., 2023) [^5] https://huggingface.co/EleutherAI/pythia-1.4b, BLOOM-1B1 (Le Scao et al., 2022) [^6] https://huggingface.co/bigscience/bloom-1b1, and TinyLlama_v1.1 (Zhang et al., 2024) [^7] https://huggingface.co/TinyLlama/TinyLlama_v1.1. The last linear layer which maps the embeddings to tokens is replaced with a linear classification head with two outputs to adapt language models to the classification setting.
For slot filling, Llama-2-7B (Touvron et al., 2023) [^8] https://huggingface.co/meta-llama/Llama-2-7b-hf was used as the strong model which yielded better performance than Qwen-7B in this task. The same set of weak models was used as the classification task. As before, both weak and strong models are finetuned with all model parameters.
@@
Baselines. The proposed Bayesian WeakS-To-Strong method is compared to the following two baselines.
@@
Joint Decoding. Joint Decoding is used as an additional baseline for text generation tasks. In contrast to the Naive Multi-Weak scheme where each weak model provides a weak target sequence, Joint Decoding employs multiple weak models to collaboratively determine one single target, as illustrated in Figure 1(b). Specifically, we perform Joint Decoding in a re-ranking fashion. For each weak model, the top $M$ sequences are generated by beam search in decoding. The sequences from the $N$ weak models are gathered to form a list of $M\times N$ sequences. Then each sequence is scored in the same way in Section 4.2. The sequence with the highest final score is used as the target sequence for strong model training. Unless otherwise mentioned, $M=5$ was used in the experiments. Each weak-to-strong experiment was run three times with different random seeds. The average and standard deviation of the results were reported. More implementation details can be found in AppendixC.
@@ 5.3 Evaluation Metrics
The classification task is evaluated by accuracy, and the SLU-F1 (Bastianelli et al., 2020) is used for slot filling, which combines both word-level and character-level F1 scores to give partial credit to non-exact match predictions. Performance gap recovered (PGR) (Burns et al., 2023) is used to measure the performance gap recovered with weak supervision, which is defined as follows:
$\text{PGR}:=\frac{\mathcal{P}-\mathcal{P}_{w}}{\mathcal{P}_{s}-\mathcal{P}_{w}}$ (13)
where $\mathcal{P}$ is Weak-to-Strong performance, $\mathcal{P}_{s}$ strong performance and $\mathcal{P}_{w}$ weak performance.
@@ 6.1 Text classification
The Naive Multi-Weak approach and Bayesian Multi-Weak using EDL were applied to the classification task for WeakS-to-Strong. Experiments with three weak models (GPT2-Large, OPT-1.3B and Pythia-1.4B) and all five weak models were conducted. With three weak models, the Naive Multi-Weak method got an average PGR of 0.661, which is lower than the single OPT-1.3B model, and Bayesian Multi-Weak boosted PGR to 0.756. This indicates that a naive ensemble approach doesn’t necessarily outperform the best single model, especially when there is a certain weak model which does not perform well. The Bayesian approach can increase the fault tolerance with the usage of prior in this case, as it learns patterns from the entire dataset. With five weak models compared to three, the Bayesian Multi-Weak approach further increased average PGR to 0.828, which is 11% higher than the best single model and 7% higher than the naive baseline, and consistently better with all seeds. The results show the effectiveness of the Bayesian approach for distribution estimation.
@@ 6.2 Text generation
For the text generation task using slot filling, the performance of the student strong model and teacher weak model finetuned on ground-truth labels is presented in Table 3. The strong ceiling performance, which is obtained by training the strong model with ground-truth labels, is 0.748, and the highest weak performance is 0.680.
The Weak-to-Strong and WeakS-to-Strong performances are reported in Table 4. For a single weak model, the Weak-to-Strong model performance didn’t necessarily surpass the original weak performance (e.g. TinyLlama_v1.1), and the highest PGR is 0.399. With five weak models, our proposed Bayesian Multi-Weak approach achieved an average PGR of 0.668, which is 26% better than a single weak model, 6% better than the naive baseline, and consistently better across three seeds. Comparing WeakS-to-Strong performance with and without auxiliary loss, adding auxiliary loss is not effective for naive approaches, but improves the performance of Bayesian approaches. It indicates that the proposed Bayesian method, where the predictions of a strong model are applied as part of the distribution estimation along with weak label as in Eqn. (5), more effectively integrates strong model predictions into the training process, showing the effectiveness of our Bayesian approach for the model ensemble. The ablation study can be found in Appendix D. A noticeable decline has been observed for Joint Decoding when the number of weak models increases from three to five. This may result from the poor performance of the TinyLlama weak model which affects the quality of the generated weak target. Recall that in Joint Decoding, multiple weak models collaboratively determine one single target. In contrast, Naive Multi-Weak and Bayesian Multi-Weak methods still benefit from the increased number of weak models, which indicates that they are more robust against the quality of a single weak model. Additional experiments on Joint Decoding can be found in Appendix F for more analysis and insights.
Table 5: Results before and after DPO. Based on the Bayesian Multi-Weak model using five weak models (the WeakS-to-Strong-5 setting) with auxiliary loss. Average results on three seeds are reported. Note that the initial SFT models are different for different seeds.
SLU-F1 PGR
Before cDPO 0.721 $\pm$ 0.013 0.668 $\pm$ 0.166
After cDPO 0.724 $\pm$ 0.015 0.705 $\pm$ 0.187
@@ Appendix C Implementation Details
All models were trained on NVIDIA A800 GPUs using the bfloat16 data type. For the classification tasks, the Adam optimizer was used with a cosine learning rate scheduler and no warm-up period. The batch size was set to 32, with a mini-batch size of 1. The weak models were finetuned on the ground-truth labels with an initial learning rate of $5\times 10^{-5}$, while the strong models were trained with a starting learning rate of $1\times 10^{-5}$ (both on the weak labels and ground-truth labels). The Weak(S)-to-Strong training was run for two epochs.
For generation tasks, the AdamW optimizer was used with a linear learning rate scheduler, also with no warm-up. The initial learning rates were set at $4\times 10^{-5}$ for GPT2-Large and Pythia-1.4B, and $8\times 10^{-5}$ for OPT-1.3B, with a batch size of 8 (mini-batch size of 4). These models were trained for 15 epochs. The checkpoints with the lowest validation loss were selected to ensure the quality of weak labels produced by the weak models. The strong model was trained with a batch size of 2 (mini-batch size of 1) and an initial learning rate of $1\times 10^{-5}$, evaluated at the end of two epochs. In the training for strong ceiling performance, the hyperparameters were adjusted based on the validation set. For Weak-to-Strong training, in which case the ground truth is not accessible, we aligned the hyperparameter settings with those used in strong ceiling training.
@@
Appendix D Ablation Study on Probability Estimation for Weak Sequences

### v4

@@ Abstract
Advances in large language models raise the question of how alignment techniques will adapt as models become increasingly complex and humans will only be able to supervise them weakly. Weak-to-Strong mimics such a scenario where weak model supervision attempts to harness the full capabilities of a much stronger model. This work extends Weak-to-Strong to WeakS-to-Strong by exploring an ensemble of weak models which simulate the variability in human opinions. Confidence scores are estimated using a Bayesian approach to guide the WeakS-to-Strong generalization. Furthermore, we extend the application of WeakS-to-Strong from text classification tasks to text generation tasks where more advanced strategies are investigated for supervision. Moreover, direct preference optimization is applied to advance the student model’s preference learning, beyond the basic learning framework of teacher forcing. Results demonstrate the effectiveness of the proposed approach for the reliability of a strong student model, showing potential for superalignment.  [^1] Supported by Shanghai Artificial Intelligence Laboratory.  [^2] Code is available in https://github.com/cuiziyun/BayesianWS2S.
@@ 3.3 Bayesian WeakS-to-Strong
For a given input $\bm{x}$, consider a weak label from the $i\text{th}$ weak model $\bm{y}_{w}^{(i)}$, which is a one-hot vector with $y_{w}^{(i,k)}$ being one if it belongs to class $k$ and zero otherwise. $\bm{y}_{w}^{(i)}$ is sampled from a categorical distribution of weak labels $\operatorname{Cat}(\bm{\pi})$, where each component ${\pi}_{k}$ corresponds to the probability assignment over the possible classes $\bm{y}_{w}^{(i)}\sim\mathrm{P}(\bm{y}|\bm{\pi})=\operatorname{Cat}(\bm{\pi}).$ EDL places a Dirichlet prior over the categorical distribution representing the probability of each categorical probability assignment, hence modelling second-order probability $\bm{\pi}\sim\mathrm{p}(\bm{\pi}|\bm{\alpha})$, where $\bm{\alpha}$ is the hyperparameter of the Dirichlet prior. The strong model $\bm{f_{\Lambda}}$ is trained to predict $\bm{\alpha}$ for each input by minimizing the negative log-likelihood of sampling $\bm{y}_{w}^{(i)}$ given the predicted Dirichlet prior:
@@
where $K$ is the number of classes, $\alpha_{0}=\sum_{k=1}^{K}\alpha_{k}$ is the Dirichlet strength, and $y_{w}^{(i,k)}$ is the $k^{th}$ value of label $\bm{y}_{w}^{(i)}$. When a sample is not correctly classified, it is expected that the prior should approach non-informative prior for this sample. Following Sensoy et al. (2018), a regularization term $\mathcal{L}_{\text{REG}}^{(i)}$ (KL-divergence between the misleading prediction and non-informative distribution, see Appendix H for details) is added to penalize incorrect predictions and calibrate uncertainty estimation, resulting in the final EDL loss $\mathcal{L}_{\text{EDL}}^{(i)}=\mathcal{L}_{\text{NLL}}^{(i)}+\lambda_{\text{EDL}}\mathcal{L}_{\text{REG}}^{(i)}$ where $\lambda_{\text{EDL}}$ is the coefficient.
@@
$\mathcal{L}_{\text{EDL}}(\bm{f_{\Lambda}}(\bm{x}),\{\bm{y}_{w}^{(i)}\}_{i=1}^{N})=\sum_{i=1}^{N}\lambda_{i}\sum_{k=1}^{K}p_{k}^{(i)}\mathcal{L}_{\text{EDL}}^{(i)}(\bm{f_{\Lambda}}(\bm{x}),\hat{\bm{y}}_{w}^{(i,k)})$ (4)
where $\hat{\bm{y}}_{w}^{(i,k)}$ is the predicted result, i.e., one-hot vector for class $k$, and $\lambda_{i}$s are hyperparameters set to the same values as used in the Naive Multi-Weak approach. As a result, the auxiliary confidence loss described in Eqn. (1) is adapted for Bayesian WeakS-to-Strong as follows:
@@ 4.1 Probability estimation for weak sequence labels
$\log\mathrm{P}(s_{1})=\frac{e^{-C_{s}(s_{1})}}{e^{-C_{s}(s_{1})}+e^{-C_{s}(s_{2})}}\log\mathrm{P}(W),\quad\log\mathrm{P}(s_{2})=\frac{e^{-C_{s}(s_{2})}}{e^{-C_{s}(s_{1})}+e^{-C_{s}(s_{2})}}\log\mathrm{P}(W),$ (7)
@@ 4.2 DPO for sequence generation optimization
where $\bm{y}_{s}(m)$ is the $m$th output sequence generated by the strong model, $\mathrm{P}(\bm{y}_{s}(m)|\bm{\theta}_{i})$ is computed using the $i$th weak model with model parameters $\bm{\theta}_{i}$, and $\lambda_{i}$ is the weight assigned to the $i$th weak model. The dispreferred sequence can be computed similarly by $\bm{y}_{r}=\operatorname*{arg\,min}\mathrm{P}(\bm{y}_{s}(m))$.
@@ 5.2 Models and baselines
Models. For classification, the Qwen-7B (Bai et al., 2023) model was applied as the strong student model. Five models were used as weak teachers: GPT2-Large (Radford et al., 2019), OPT-1.3B (Zhang et al., 2022), Pythia-1.4B (Biderman et al., 2023), BLOOM-1B1 (Le Scao et al., 2022), and TinyLlama_v1.1 (Zhang et al., 2024). The last linear layer which maps the embeddings to tokens is replaced with a linear classification head with two outputs to adapt language models to the classification setting.
For slot filling, Llama-2-7B (Touvron et al., 2023) was used as the strong model which yielded better performance than Qwen-7B in this task. The same set of weak models was used as the classification task. As before, both weak and strong models are finetuned with all model parameters.
@@
Baselines. The proposed Bayesian WeakS-to-Strong method is compared to the following three baselines.
@@
FlyingSquid. FlyingSquid (Fu et al., 2020) is a method for weak supervision, estimating the accuracies and correlations among multiple noisy label functions (different weak labels in our case) without ground-truth data. Latent variable probabilistic graphical models are used to model these dependencies, with weak labels as observed variables and unobserved ground-truth labels as hidden variables. Since FlyingSquid is designed for binary classification, this baseline is only used for classification. Through this method, we get a label model with multiple weak labels and obtain the probability for the positive category, which is then used as a soft label in strong model training.
Joint Decoding. Joint Decoding is used as an additional baseline for text generation tasks, which is specifically designed for multiple weak model generations. In contrast to the Naive Multi-Weak scheme where each weak model provides a weak target sequence, Joint Decoding employs multiple weak models to collaboratively determine one single target, as illustrated in Figure 1(b). Specifically, we perform Joint Decoding in a re-ranking fashion. For each weak model, the top $M$ sequences are generated by beam search in decoding. The sequences from the $N$ weak models are gathered to form a list of $M\times N$ sequences. Then each sequence is scored in the same way in Section 4.2. The sequence with the highest final score is used as the target sequence for strong model training. Unless otherwise mentioned, $M=5$ was used in the experiments.
Each weak-to-strong experiment was run three times with different random seeds. The average and standard deviation were reported. More implementation details can be found in Appendix C.
@@ 5.3 Evaluation Metrics
The classification task is evaluated by accuracy, and the SLU-F1 (Bastianelli et al., 2020) is used for slot filling, which combines both word-level and character-level F1 scores to give partial credit to non-exact match predictions. Performance gap recovered (PGR) (Burns et al., 2023) is used to measure the performance gap recovered with weak supervision, which is defined as $(\mathcal{P}-\mathcal{P}_{w})/(\mathcal{P}_{s}-\mathcal{P}_{w})$ where $\mathcal{P}$ is Weak-to-Strong performance, $\mathcal{P}_{s}$ strong performance and $\mathcal{P}_{w}$ weak performance. For multiple weak model cases, the average of PGRs for each weak model is treated as the final result.
@@ 6.1 Text classification
The Naive Multi-Weak baseline, FlyingSquid approach and Bayesian Multi-Weak using EDL were applied to the classification task for WeakS-to-Strong. Experiments with three weak models (GPT2-Large, OPT-1.3B and Pythia-1.4B) and all five weak models were conducted. With three weak models, the Naive Multi-Weak method got an average PGR of 0.661, slightly outperforming FlyingSquid method but still lower than the single OPT-1.3B model, and Bayesian Multi-Weak boosted PGR to 0.756. This indicates that a naive ensemble approach doesn’t necessarily outperform the best single model, especially when there is a certain weak model which does not perform well. The Bayesian approach can increase the fault tolerance with the usage of prior in this case, as it learns patterns from the entire dataset. With five weak models compared to three, the Bayesian Multi-Weak approach further increased average PGR to 0.828, which is 16% relatively higher than the best single model and 8% relatively higher than baselines, and consistently better with all seeds. The results show the effectiveness of the Bayesian approach for distribution estimation.
@@ 6.2 Text generation
For the text generation task on slot filling, the performance of the student strong model and teacher weak model finetuned on ground-truth labels is presented in Table 3. The strong ceiling performance is 0.748, and the highest weak performance is 0.680.
The Weak(S)-to-Strong performances are reported in Table 4. For a single weak model, the Weak-to-Strong model performance didn’t necessarily surpass the original weak performance (e.g. TinyLlama_v1.1), and the highest PGR is 0.399. With five weak models, our proposed Bayesian Multi-Weak approach achieved an average PGR of 0.668, which is 26% better than a single weak model, 6% better than the naive baseline, and consistently better across three seeds. Comparing WeakS-to-Strong performance with and without auxiliary loss, adding auxiliary loss is not effective for naive approaches, but improves the performance of Bayesian approaches. It indicates that the proposed Bayesian method, where the predictions of a strong model are applied as part of the distribution estimation as in Eqn. (5), more effectively integrates strong model predictions into the training process, showing the effectiveness of our Bayesian approach for the model ensemble. The ablation study can be found in Appendix E. A noticeable decline has been observed for Joint Decoding when the number of weak models increases from three to five. This may result from the poor performance of the TinyLlama weak model which affects the quality of the generated weak target. Recall that in Joint Decoding, multiple weak models collaboratively determine one single target. In contrast, Naive Multi-Weak and Bayesian Multi-Weak methods still benefit from the increased number of weak models, which indicates that they are more robust against the quality of a single weak model. Additional experiments on Joint Decoding can be found in Appendix G for more analysis and insights.
Table 5: Results before and after DPO. Based on the Bayesian Multi-Weak model using five weak models (the WeakS-to-Strong-5 setting) with auxiliary loss. Average results on three seeds are reported. Note that for different seeds, the initial SFT models are different.
seed-0 seed-1 seed-2 Average
SLU-F1 PGR SLU-F1 PGR SLU-F1 PGR SLU-F1 PGR
Before cDPO 0.706 0.477 0.730 0.776 0.728 0.751 0.721 $\pm$ 0.013 0.668 $\pm$ 0.166
After cDPO 0.707 0.490 0.733 0.813 0.733 0.813 0.724 $\pm$ 0.015 0.705 $\pm$ 0.187
@@ Appendix C Implementation Details
All models were trained on NVIDIA A800 GPUs using the bfloat16 data type. For the classification tasks, the Adam optimizer was used with a cosine learning rate scheduler and no warm-up period. The batch size was set to 32, with a mini-batch size of 1. The weak models were finetuned on the ground-truth labels with an initial learning rate of $5\times 10^{-5}$, while the strong models were trained with a starting learning rate of $1\times 10^{-5}$ (both on the weak labels and ground-truth labels). The Weak(S)-to-Strong training was run for two epochs. The weights for different weak models ($\lambda$) was set to average values to simplify the classification experiments (refer to Appendix F for a comparison between average and fixed weights).
For generation tasks, the AdamW optimizer was used with a linear learning rate scheduler, also with no warm-up. The initial learning rates were set at $4\times 10^{-5}$ for GPT2-Large and Pythia-1.4B, and $8\times 10^{-5}$ for OPT-1.3B, with a batch size of 8 (mini-batch size of 4). These models were trained for 15 epochs. The checkpoints with the lowest validation loss were selected to ensure the quality of weak labels produced by the weak models. The strong model was trained with a batch size of 2 (mini-batch size of 1) and an initial learning rate of $1\times 10^{-5}$, evaluated at the end of two epochs. In the training for strong ceiling performance, the hyperparameters were adjusted based on the validation set. For Weak-to-Strong training, in which case the ground truth is not accessible, we aligned the hyperparameter settings with those used in strong ceiling training. The weights $\lambda$ in Eqn. 4 were set to (0.1, 0.3, 0.2, 0.3, 0.1) for GPT2-Large, OPT-1.3B, Pythia-1.4B, BLOOM-1B1 and TinyLlama_v1.1 respectively.
@@
Appendix D Experiments on an Additional Dataset
To verify the generalization of our method, additional experiments were conducted on another dataset, CosmosQA (Huang et al., 2019), for the classification task. CosmosQA is a large-scale dataset of problems that require commonsense-based reading comprehension. Like the preprocess on the SciQ dataset, 5k data samples were extracted for training weak models, another 5k samples for strong models, and 1k samples for testing. This dataset was also reformatted into a binary classification format (i.e., determining correctness). Experiments with three weak models were conducted. The weak model performance and strong performance are listed in Table 6 and Weak(S)-to-Strong performance in Table 7.
Table 6: Performance of single model on text classification task on CosmosQA dataset. Trained by ground-truth labels.
Pre-trained model # Param Accuracy
Strong Model (ceiling) Qwen-7B 7.7B 0.847
Weak Model GPT2-Large 0.8B 0.642
OPT-1.3B 1.3B 0.642
Pythia-1.4B 1.4B 0.654
Table 7: Weak(S)-to-Strong performance on text classification task for with and without auxiliary loss on CosmosQA dataset. Experiments with 3 weak models were conducted. Each experiment was run with three different random seeds. The best results are shown in bold.
w/o aux loss w/ aux loss
Accuracy PGR Accuracy PGR
Weak-to-Strong GPT2-Large 0.652 $\pm$ 0.007 0.077 $\pm$ 0.032 0.651 $\pm$ 0.007 0.073 $\pm$ 0.032
OPT-1.3B 0.687 $\pm$ 0.005 0.218 $\pm$ 0.022 0.704 $\pm$ 0.009 0.304 $\pm$ 0.044
Pythia-1.4B 0.685 $\pm$ 0.010 0.162 $\pm$ 0.052 0.731 $\pm$ 0.002 0.399 $\pm$ 0.009
WeakS-to-Strong-3 Naive Multi-Weak 0.691 $\pm$ 0.009 0.232 $\pm$ 0.043 0.698 $\pm$ 0.007 0.267 $\pm$ 0.034
FlyingSquid 0.699 $\pm$ 0.004 0.272 $\pm$ 0.020 0.706 $\pm$ 0.004 0.303 $\pm$ 0.030
Bayesian Mutli-Weak 0.694 $\pm$ 0.007 0.247 $\pm$ 0.033 0.760 $\pm$ 0.005 0.571 $\pm$ 0.025
Table 6 shows that, on the CosmosQA dataset, both the weak and strong models perform worse compared to that on the SciQ dataset, with the weak model’s performance dropping more. In the Weak(S)-to-Strong experiments, neither multi-weak baseline outperformed the best single model (Pythia-1.4B with auxiliary loss), while our Bayesian approach significantly exceeded it, showing the effectiveness of our method, especially when weak model performance varies.
Appendix E Ablation Study on Probability Estimation for Weak Sequences